A person who is 6 feet tall walks away from a 50-foot silo toward the tip of the silo's shadow. At a distance of 32 feet feet from the silo, the person's shadow begins to emerge beyond the silo's shadow. How much farther must the person walk to be completely out of the silo's shadow?
Answer by scott8148(6628)

6 / x = 50 / (32 + x)

192 + 6x = 50x

192 = 44x ___ 43 / 11 = x